OpenStudy (anonymous):

-5/6x -5/9 + 2/3x= -28/45 / Another clearing fraction problem

OpenStudy (saifoo.khan):

-1x/6 = -1/15

jhonyy9 (jhonyy9):

check the denominator common and you will get 90x - so for this the first term you need to multiplied with 15 the second with 10x the 3th with 30 and the 4th with 2x ,and in this way you will get -75 -50x +60=-56x -15=-6x 15=6x x=15/6 =5/3